OpenStudy (anonymous):

for the given functions f and g, find the indicated composition. f(x)=17x^2-10, g(x)=19x-10 a) (fog)(x) b) (gof)(-3)

OpenStudy (anonymous):

can someone please show the steps for this question?

OpenStudy (anonymous):

a) f(g(x))=17(19x-10)^2-10 =17(361x²-380x+100)-10 =6137x²-6460x+1690 not shure =/

OpenStudy (anonymous):

you put the 'g' values for 'x' in the place of 'x' in 'f'

OpenStudy (anonymous):

what about the (-3)?

OpenStudy (anonymous):

b) g(f(x))=19(17x²-10)-10 if x=3 => 19(143)-10 =2717-10 =1707

OpenStudy (anonymous):

for x= -3, sorry. the answer is still the same

OpenStudy (anonymous):

2707, sorry

OpenStudy (anonymous):

thanks alot, but how can I get a final answer on the first part?

OpenStudy (anonymous):

can i simply imput that into a calculator?

OpenStudy (anonymous):

wow, I get 37,657,999 is that correct? It seems like to large of a number.

OpenStudy (anonymous):

only if you have a number to 'x'. in this case, the aswer is the function.

OpenStudy (anonymous):

there is no number because there isn't a value for x.

OpenStudy (anonymous):

forgot to use the x sorry

OpenStudy (anonymous):

Are you sure that is the correct final equation?

OpenStudy (anonymous):

not SHURE, but shure ^^"

OpenStudy (anonymous):

did u understand the principle? I think it's the most important thing =D

OpenStudy (anonymous):

yes, but I not sure why the "x" shows up in the first part, but not the second part

OpenStudy (anonymous):

because in the second part, it gives you the '-3'. when it should says f(x), it says f(-3). So you put the number in place of 'x'.

OpenStudy (anonymous):

got it, thank you very much. Do you have time for another(but different )problem? If you want me to post to everyone, thats o.k too

OpenStudy (anonymous):

better post to everyone ^^" it's 11 PM here. going to sleep ;D

OpenStudy (anonymous):

no problem, thanks again

OpenStudy (anonymous):

u're welcome